# Half-metallicity and wandering axis ferromagnetism in Mn substituted Fe2TiSn

### v1

@@ III Results and Discussion
For small values of applied magnetic field, there is a noticeable contrast between the M(T) plots recorded during the zero field cooled (ZFC) and field cooled (FC) protocols. This bifurcation between ZFC/ FC disappears when a relatively strong magnetic field of 0.1 T is applied (Fig.8), indicating the presence of weak anisotropy in the material. In the ZFC process, the material’s anisotropy determines the amount of magnetization below TC and may prevent a relatively small applied field from aligning the moments in the direction of the field. When the same material is cooled through TC while being exposed to a magnetic field during the FC process, the moments are locked into the direction of the applied magnetic field. As a result, for small applied fields, there is a noticeable contrast between the ZFC and FC curves, and this bifurcation disappears when a slightly stronger magnetic field is applied.
As seen from the electronic structure calculations, the substitution of Mn in Fe2TiSn enhances the magnetic properties by introducing strong local moments at Mn sites. Accordingly, the isothermal magnetization M(H) of all the Mn–compositions measured at 5 K shows a sizable increase in the magnetic moment values at a small applied field (see Fig.9(a)). The magnetic moment per formula unit extracted from extrapolating the near saturation M(H) data, is found to exactly follow the Slater-Pauling rule (see Fig. 9(b)), reflecting the half-metallic state nature of Fe2Ti1-xMnxSn compositions. However, it may be noted that the M(H) curves do not reach complete saturation, nor yield a good fit to the usual modified Langevin function, , where = H/kBT, MS represent the saturation magnetization, is the average magnetic moment and represents the paramagnetic susceptibility. Thus ruling out any magnetic cluster-like evolution in the prepared compositions.
@@
Alternatively, from the measurements, it appears that the Mn-substituted Fe2TiSn compositions display a long-range ferromagnetic order below TC, which exhibits a substantial dependency on the strength of the applied magnetic field with respect to the random anisotropy. Further, the (T) data shows the Kondo-like behavior i.e. the presence of local Mn moment screened by conduction electrons, a weak random anisotropy model[34] could be applicable here. Such a model has previously been extensively used to explain the short-range FM correlations in Huesler alloys[35, 36, 37]. This model incorporates both the experimentally applied magnetic field and the ground state configuration of magnetic materials with random anisotropy for a wide range of anisotropy strengths. Three distinct regimes can be identified for the weak anisotropy situation based on the magnitude of the applied magnetic field, H, and the parameter HS (= H4r/H3ex), where Hr and Hex are anisotropic and exchange fields, respectively. For H HS, one obtains a correlated spin glass phase with a high magnetic susceptibility for low fields. When HS H Hex, a ferromagnet with a wandering axis (FMWA) results from a rough alignment of the spins in the presence of a weak magnetic field. The tipping angle of the magnetization vector with respect to the applied field changes throughout the system due to the non-collinear magnetic structure. In this regime, the magnetization approaches saturation as,
@@
where Ms is saturation magnetization, and Hc is the field due to the coherent portion of the anisotropy. In the third regime where H HS, the correlation length of the spins reduces as the field strength increases, whereas the tipping angle is completely uncorrelated amongst sites. Except for a slight tipping angle (less than in the case of FMWA) brought on by random anisotropy, all spins in this regime are almost aligned with the field. In such cases, isothermal magnetization gradually approaches saturation as,
@@
We have fitted equation(1) with an additional paramagnetic contribution,H[38], to the M(H) curve at 5 K of = 0.25, as shown in Fig.10. It shows a satisfactory fit for a wide range of applied fields from 0.4 to 5 T. The value of Hs derived after the fitting ranges from 0.303 T for = 0.05 to 0.021 T for = 0.30, which is lower than the maximum applied field, following the condition for the system to fall in the second regime. In contrast, the fitting of equation(2) does not yield a good fit to the data (see the dashed line in Fig.10 and the value of Hs derived after the fitting is much higher than the maximum applied field despite using the prefactor equal to 1 instead of 1/15[36].
@@
Thus, the analysis of ab initio DOS calculation, resistivity, and magnetic data suggests that Mn substitution in Fe2TiSn introduces localized moments. These local moments exhibit strong coupling with the conduction electron, which screens the interaction between two Mn atoms. As the results of this screening, the strength of the magnetic interaction decreases with increasing Mn-Mn distance, and typical ferromagnetic order is absent in Fe2Ti1-xMnxSn. To verify this, we have calculated the distance-dependent magnetic exchange interaction (JMn-Mn) between two Mn atoms. Estimation of distance-dependent Mn-Mn magnetic interaction with smaller unit cells (considered for 25 substitution) is a difficult task. This is because, at large Mn-Mn separation in a small unit cell, contributions from the periodic images of Mn atoms at relatively shorter distances will start contributing to the magnetic energy. To avoid this problem, one has to consider a large supercell. We consider a 222 supercell and replace two Ti with Mn atoms. The distance-dependent Mn-Mn magnetic interaction (JMn-Mn = EAFM-EFM) is then estimated as explained in the ”Experimental and Computational details” section. Here, we take advantage of the fact that the magnetic moments on Fe and Ti atoms are negligibly small, and hence their insignificant direct contributions to JMn-Mn can safely be ignored. The plot of separation dependent JMn-Mn is shown in Fig. 11. One can observe that the effective Mn-Mn interaction is ferromagnetic, consistent with the experimental findings. This ferromagnetic interaction exponentially decreases with the Mn-Mn separation and nearly zeroed at the separation of 8.51Å.
@@
Hence, according to the results of magnetic measurements, the magnetic state of Fe2Ti1-xMnxSn changes with temperature as marked in the -T phase diagram, as shown in the Fig.12(a). Fe2Ti1-xMnxSn is in a typical PM state when the temperature is higher than Ts. As temperature decreases ferromagnetic correlation appears due to the formation of magnetic clusters. In the case of (region-I), clusters continue to evolve with falling temperatures and lead to a superparamagnetic ground state. No long-range magnetic order is observed in this region. The simplified schematic diagram of the magnetic state as a function of temperature is presented in Fig.12(b). Substituting higher Mn content ( ) in place of Ti in Fe2Ti1-xMnxSn significantly alters this magnetic behavior. The strong local moments of Mn suppress the influence of high-temperature FM clusters on the low-temperature magnetic ground state and result in an unconventional ferromagnetic ordering driven by random anisotropy, known as the ”ferromagnet with wandering axis” state, as represented as region-II in the phase diagram. This transition from a superparamagnetic state in Fe2TiSn to a weak random anisotropy-driven state with Mn substitution underscores the significant impact of Mn substitution on the magnetic properties of Heusler alloys. Furthermore, as confirmed from DFT calculation, electrical resistivity, and magnetic measurements, Mn substitution induces half-metallicity in the Fe2Ti1-xMnxSn, enhancing its potential for spintronic applications.
@@ IV Conclusion
In this study, structural, magnetic, and transport properties of Mn substituted Fe2TiSn have been investigated through room temperature XRD, temperature and field-dependent magnetization, and magnetotransport measurements. The room temperature powder X-ray diffraction patterns of all the Fe2Ti1-xMnxSn compositions are indexed to Fmm symmetry with an ordered L21 crystal structure. The lattice constants decrease systematically with rising Mn content. The temperature variation of electrical resistivity Fe2Ti1-xMnxSn shows an overall increase in the magnitude with increasing Mn content. For 0.25, (T) shows a negative temperature coefficient of resistance, followed by an upturn at T 20 K. The low-temperature upturn is due to the combined effect of electron-electron correlation and strong Kondo-like spin scattering. The (T) above 100 K is dominated by the electron-phonon scattering, with a small contribution from the electron-magnon scattering as well, which indicates the presence of spin-polarized density of state in these compositions. Electrical transport along with the density of states calculation confirms the formation of highly spin-polarized electronic structures and the presence of half-metallicity on Mn substituted Fe2TiSn.
Magnetization recorded as a function of temperature exhibits signatures akin to magnetic ordering and shifts to higher temperatures with increasing Mn concentration. Such a trend indicates the presence of a strong local magnetic moment of Mn atoms. However, the same M(T) curves, upon application of magnetic fields of higher magnitude display features that hint towards the presence of anisotropy in the system. Besides, a step-like feature is recorded at T 250 K in all the M(T) curves. Field-dependent magnetization, M(H), recorded at fixed temperatures shows the presence of ferromagnetic correlations even at 300 K. The low-temperature M(H) suits a random anisotropy model suggesting that a ferromagnet with a wandering axis can be an apt description of the magnetic ground state of the Fe2Ti1-xMnxSn compositions.
V Acknowledgement
K.M. acknowledges DST-INSPIRE (File: DST/INSPIRE/03/2021/001384/IF190777), New Delhi, for providing the research fellowship. KM and PB thank Dr. Mukul Gupta, UGC-DAE CSR, Indore, for extending the X-ray diffraction facility. Acknowledgment is also due to Prof. J. G. Lin, Center for Condensed Matter Science, National Taiwan University, Taipei, Taiwan, for providing access to the magnetic measurement facility. We also thank Prof. Pratap Raychaudhuri and Ganesh Jangam for their help in magnetic measurement at the Tata Institute of Fundamental Research, Mumbai.

### v2

@@ III Results and Discussion
For small values of applied magnetic field, there is a noticeable contrast between the M(T) plots recorded during the zero field cooled (ZFC) and field cooled (FC) protocols. This bifurcation between ZFC/ FC disappears when a relatively strong magnetic field of 0.1 T is applied (Fig.8), indicating the presence of weak anisotropy in the material. In the ZFC process, the material’s anisotropy determines the amount of magnetization below TC and may prevent a relatively small applied field from aligning the moments in the direction of the field. When the same material is cooled through TC while being exposed to a magnetic field during the FC process, the moments are locked into the direction of the applied magnetic field. As a result, there is a noticeable contrast between the ZFC and FC curves for small applied fields, and this bifurcation disappears when a slightly stronger magnetic field is applied.
Fig.9(a-b) shows the variation of isothermal magnetization () as a function of the applied magnetic field () for all the compositions measured at 300 K and 5 K, respectively. For = 0.05, the curve at 300 K displays a linear behavior (Fig.9(a)). Further increase in , a sharp rise in at lower fields is observed followed by a linear component at higher fields. These results indicate the presence of weak ferromagnetic clusters at 300 K. Consequently, the curve measured at = 5 K shows characteristics of a ferromagnetic ordering with an increase in the magnetic moment values with (see Fig.9(b)). The saturation magnetic moment per formula unit is found to follow the Slater-Pauling rule (Fig. 9(c)), reflecting possibly the spin-polarized band structure of Fe2Ti1-xMnxSn compositions. However, it may be noted that the curves do not reach complete saturation till = 5 T. The presence of ferromagnetic clusters at = 300 K suggests that the low-temperature magnetization behavior could be due to the evolution of these clusters. To interpret this, we have fitted the curve using the modified Langevin function, expressed as , where = H/kBT, MS represents the saturation magnetization, is the average magnetic moment and represents the paramagnetic susceptibility. The lack of fit of this model to the data (not shown here) rules out any cluster-like evolution in the prepared compositions.
@@
Alternatively, from the measurements, it appears that the Mn-substituted Fe2TiSn compositions display a long-range ferromagnetic order below , which exhibits a substantial dependency on the strength of the applied magnetic field with respect to the random anisotropy. Further, the data shows the Kondo-like behavior, i.e., the presence of local Mn moment screened by conduction electrons, a weak random anisotropy model[43] could be applicable here. Such a model has previously been extensively used to explain the short-range FM correlations in Heusler alloys[44, 45, 46]. This model incorporates both the experimentally applied magnetic field and the ground state configuration of magnetic materials with random anisotropy for a wide range of anisotropy strengths. Three distinct regimes can be identified for the weak anisotropy situation based on the magnitude of the applied magnetic field, H, and the parameter (= r/ex), where and are anisotropic and exchange fields, respectively. For , one obtains a correlated spin glass phase with a high magnetic susceptibility for low fields. When , a ferromagnet with a wandering axis (FMWA) results from a rough alignment of the spins in the presence of a weak magnetic field. The tipping angle of the magnetization vector, with respect to the applied magnetic field, changes throughout the system due to the non-collinear magnetic structure. In this regime, the magnetization approaches saturation as,
@@
where is saturation magnetization, and is the field due to the coherent portion of the anisotropy. In the third regime where , the correlation length of the spins reduces as the field strength increases, whereas the tipping angle is completely uncorrelated amongst sites. Except for a slight tipping angle (less than in the case of FMWA) brought on by random anisotropy, all spins in this regime are almost aligned with the field. In such cases, isothermal magnetization gradually approaches saturation as,
@@
We have fitted equation(1) with an additional paramagnetic contribution,H[47], to the M(H) curve at 5 K of = 0.25, as shown by the solid line in Fig.10. It shows a satisfactory fit for a wide range of applied fields from 0.4 to 5 T. The value of Hs derived after the fitting ranges from 0.303 T for = 0.05 to 0.021 T for = 0.30, which is lower than the maximum applied field, following the condition for the system to fall in the second regime. In contrast, equation(2) does not yield a good fit to the data (see the dashed line in Fig.10), and the value of derived after the fitting is much higher than the maximum applied field despite using the prefactor equal to 1 instead of 1/15[45].
Random anisotropy in cubic systems can occur from substitutional impurities and disorder[48]. In Ref.[47], P.M. Gehring et al. proposes that the random anisotropy in cubic polycrystalline Dy1-xYxAl2 could originate from lowering of local lattice symmetry, driven by lattice mismatch, difference in charge screening and spin-orbit interactions[47]. In certain cubic Heusler systems, such random anisotropy is found to influence its low-temperature magnetic properties [44]. To unravel the reason that the random anisotropy in our cubic Fe2Ti1-xMnxSn originates from local lattice distortion, we undertake measurement of the high-resolution synchrotron-based EXAFS spectra at Fe and Ti K-edges for = 0.25. The EXAFS data is extracted in the reciprocal space, , and Fourier transformed to real space, , where the magnitude and real part of reflects the radial distribution function around the absorbing atoms. The local structural refinement is carried out by varying the atomic distances and thermal mean square factor () until a good fit is obtained in the range of 1.8–3.0 Å in the space and 3 to 14 Å-1 in space. Based on the information extracted from the XRD profile of the cubic L21 structure and lattice constant of 6.059 Å , the local atomic arrangement around the Fe atom has 4 Ti and 4 Sn atoms at an equal distance of 2.623 Å, and 6 Fe atoms at 3.029 Å. Similarly, for the Ti atom, there are 8 Fe atoms at 2.623 Å and 6 Sn atoms at 3.029 Å. Given its cubic symmetry, all these bond distances can be represented as a multiple of lattice constant, and, in principle, only varying the lattice constant should give a good fit. However, this model does not fit the experimental data satisfactorily. Whereas a better fit is achieved by independently varying the individual bond distances, as shown in Fig.11, and the obtained fitting parameters are listed in Table-3. The notable aspect is the discrepancy in the Fe–Ti bond distance when viewed separately from the Fe and Ti edges. From the Ti-edge, the Ti–Fe bond distance comes out to be 2.613 Å, consistent with the value derived from the lattice constant, whereas the same bond distance, when viewed from the Fe edge, yields a value of 2.596 Å. Furthermore, the Fe–Sn bond distance observed from the Fe-edge is 2.621 Å, which aligns well with the lattice constant. This observation contradicts the expectations of the cubic symmetry where Fe–Ti and Fe–Sn bond distances should be equidistant from the Fe atom. This discrepancy in the Fe–Ti correlation comes from the fact that Mn replaces some Ti atoms in Fe2Ti1-xMnxSn compositions, and the bond distance involving the Ti site represents an average distance of Fe–Ti and Fe–Mn. This further indicates that in the Fe–(Ti/Mn) correlation, where Mn occupies the vacant Ti sites, the local lattice environment is distorted, leading to variations in the bond distance. These observations confirm that it is the lowering of local symmetry which causes the random anisotropy in Fe2Ti1-xMnxSn compositions.
Bond type | CN | R(Å) | 2(Å2) | |
Ti K-edge | ||||
Ti-Fe | 8 | 2.623 | 2.613(5) | 0.00376(38) |
Ti-Sn | 6 | 3.029 | 3.023(9) | 0.00543(76) |
Fe K-edge | ||||
Fe-Sn | 4 | 2.623 | 2.621(1) | 0.00307(15) |
Fe-Ti/Mn | 4 | 2.623 | 2.596(3) | 0.00568(45) |
Fe-Fe | 6 | 3.029 | 3.008(4) | 0.00767(43) |
@@
Thus, the analysis of ab initio DOS calculation, resistivity, and magnetic data suggest that Mn substitution in Fe2TiSn introduces localized moments. These local moments exhibit strong coupling with the conduction electron, which screens the interaction between two Mn atoms. As the results of this screening, the strength of the magnetic interaction decreases with increasing Mn-Mn distance, and typical ferromagnetic order is absent in Fe2Ti1-xMnxSn. We have calculated the distance-dependent magnetic exchange interaction (JMn-Mn) between two Mn atoms to verify this. Estimating distance-dependent Mn-Mn magnetic interaction with smaller unit cells (considered for 25 substitution) is difficult. This is because, at large Mn-Mn separation in a small unit cell, contributions from the periodic images of Mn atoms at relatively shorter distances will start contributing to the magnetic energy. To avoid this problem, one has to consider a large supercell. We consider a 222 supercell and replace two Ti with Mn atoms. The distance-dependent Mn-Mn magnetic interaction (JMn-Mn = EAFM-EFM) is then estimated as explained in the ”Experimental and Computational details” section. Here, we take advantage of the fact that the magnetic moments on Fe and Ti atoms are negligibly small, and hence, their insignificant direct contributions to JMn-Mn can safely be ignored. The plot of separation dependent JMn-Mn is shown in Fig. 12. One can observe that the effective Mn-Mn interaction is ferromagnetic, consistent with the experimental findings. This ferromagnetic interaction exponentially decreases with the Mn-Mn separation and becomes negligible at the separation of 8.51Å.
@@
Hence, according to the results of magnetic measurements, the magnetic state of Fe2Ti1-xMnxSn changes with temperature as marked in the -T phase diagram, as shown in the Fig.13(a). Fe2Ti1-xMnxSn is in a typical PM state when the temperature is higher than Ts. As temperature decreases, ferromagnetic correlation appears due to the formation of magnetic clusters. In the case of (region-I), clusters continue to evolve with falling temperatures and lead to a superparamagnetic ground state, and no long-range magnetic order is observed in this region. The simplified schematic diagram of the magnetic state as a function of temperature is presented in Fig.13(b-c). Substituting higher Mn content ( ) in place of Ti in Fe2Ti1-xMnxSn significantly alters this magnetic behavior. The strong local moments of Mn suppress the influence of high-temperature FM clusters on the low-temperature magnetic ground state and result in an unconventional ferromagnetic ordering driven by random anisotropy, known as the ”ferromagnet with wandering axis” state, as represented as region-II in the phase diagram. This transition from a superparamagnetic state in Fe2TiSn to a weak random anisotropy-driven state with Mn substitution underscores the significant impact of Mn substitution on the magnetic properties of Heusler alloys. Furthermore, as confirmed from electrical resistivity and magnetic measurements and further supported through our DFT calculations, Mn substitution induces half-metallicity in the Fe2Ti1-xMnxSn, enhancing its potential for spintronic applications.
It is crucial to highlight that analysis of low-temperature resistivity and isothermal magnetization presents two seemingly divergent scenarios: the Kondo effect and ferromagnetic order and coexist in the same system. The Kondo effect is generally suppressed in a strong ferromagnet characterized by high interaction strength and ordering temperature because magnetic ordering fixes the local moment direction. However, there have been reports of simultaneous observation of Kondo scattering and ferromagnetism in heavy fermion systems with partially filled 4f or 5f shells[31, 28, 49, 50], which are theoretically modeled as a regular lattice of local moments coupled to conduction electrons, and described as underscreened Kondo lattice model[51]. Besides, this observation is not limited to the heavy fermionic system with f-spins but also observed in simple 3d electron systems. For instance, Pasupathy et al. have observed Kondo-assisted tunneling via C60 molecules in contact with ferromagnetic nickel electrodes[52]. Additionally, thin films of TiO2 and Pt-substituted MnBi have shown the coexistence of ferromagnetism and Kondo effect[53, 54]. Recently, the ferromagnetism-induced Kondo effect has been observed in magnetically doped graphene, and this doped ferromagnetic graphene behaves as half-metal[55]. Moreover, Kondo-dominated low-temperature resistivity is also observed in Fe2-xNixTiSn Heusler system[15].
@@ IV Conclusion
This study investigated structural, magnetic, and transport properties of Mn substituted Fe2TiSn through room temperature XRD, temperature and field-dependent magnetization, and magnetotransport measurements. Room temperature XRD patterns confirm the single phase formation of all the compositions with order L21 structure. The temperature variation of electrical resistivity Fe2Ti1-xMnxSn shows an overall increase in the magnitude with increasing Mn content. For 0.25, shows a negative temperature coefficient of resistance, followed by an upturn at 20 K due to Kondo-like scattering and EEI. High temperature is dominated by electron-phonon scattering, indicating spin-polarized states. Electrical transport and the density of states calculation confirm the formation of highly spin-polarized electronic structures and the presence of half-metallicity on Mn substituted Fe2TiSn. Magnetization measurements display features indicative of weak anisotropy in the system, and low-temperature behavior aligns with the random anisotropy model, suggesting that a ferromagnet with a wandering axis can be an apt description of the magnetic ground state of the Fe2Ti1-xMnxSn compositions. The local lattice distortion in the prepared compositions leads to weak random anisotropy, affecting the overall magnetic properties.
V Acknowledgment
K.M. acknowledges DST-INSPIRE (File: DST/INSPIRE/03/2021/001384/IF190777), New Delhi, for providing the research fellowship. KM and PB thank Dr. Mukul Gupta, UGC-DAE CSR, Indore, for extending the X-ray diffraction facility. Acknowledgment is also due to Prof. J. G. Lin, Center for Condensed Matter Science, National Taiwan University, Taipei, Taiwan, for providing access to the magnetic measurement facility. We also thank Prof. Pratap Raychaudhuri and Ganesh Jangam for their help in magnetic measurement at the Tata Institute of Fundamental Research, Mumbai. K.M. thanks Mr. Rajeev Joshi, UGC-DAE CSR, for helping with electrical transport measurements. Part of this research is carried out at the light source PETRA III at DESY, a member of the Helmholtz Association (HGF). KM and PB thank Dr. Edmund Welter for his assistance in using the photon beamline P65. We acknowledge the financial support from the Department of Science and Technology (Govt. of India) provided within the framework of the India@DESY collaboration).